Find the equation of the line through (9, -6) with slope -3.
Answer by stanbon(75887)

Find the equation of the line through (9, -6) with slope -3.
------
Form: y = mx+b
Substitute for y,m,x and solve for "b":
-6 = -3*9 + b
-6 = -27 + b
b = 21
------
Equation:
y = -3x+21
==============
Cheers,
Stan H.